# Quasi-likelihood inference for SDE with mixed-effects observed at high frequency

### v1

@@ 1. Introduction
Mixed-effects models play a pivotal role in various scientific domains, facilitating the precise analysis of repeated observations among individuals. Despite their utility, parameter estimation in these models presents significant technical challenges. However, their application improves our understanding of dynamic biological phenomena, making them invaluable tools in numerous fields. Moreover, there is a growing interest in employing mixed-effects models coupled with stochastic differential equations (SDE). Originating from efforts in pharmacokinetics (see , e.g. references in [12]), where the aim was to refine the modeling of intrinsic variability in drug distribution within organisms, these models have now branched out into diverse applications within life sciences. From the intricate workings of neurobiology ([11]) to the intricate dance of epidemic dynamics prediction ([21]), mixed-effects models offer invaluable insight. Tailored estimation methodologies have been developed to fully exploit the potential of these models. Although numerous algorithms are proposed in the literature (see e.g. references in [24]), addressing parameter estimation in these models from a theoretical point of view remains rare. Indeed, maximum likelihood estimation, a favored method, poses challenges due to its rare explicit form. Only in select works do the authors provide specific results. Ditlevsen and De Gaetano mainly limited to Brownian motion with random drift [10]. Then Delattre, Samson, Genon-Catalot, and Larédo focused on real diffusions with specific characteristics and Gaussian and inverse gamma random effects in the drift and in the diffusion coefficient, respectively ([5], [6], [7], [8]). Covariates and Gaussian random effects in the drift were considered in [18] in multivariate diffusion models. However, there still remains a gap in understanding the asymptotic behavior of parameter estimators in more general mixed-effects diffusion models.
@@
Furthermore, prevalent assumptions in mixed-effects model literature, such as random effects following Gaussian distributions, may not always hold in practice. Recent advancements propose extensions, introducing distributions like the generalized hyperbolic distribution to better accommodate variability (see e.g. [2] and [16]).
@@
Our work revisits the notion of utilizing normal variance-mean mixture type random effects in broader mixed-effects models defined by stochastic differential equations. This allows for the exploration of a wider class of mixed-effects diffusion models compared to the previous literature. We also consider that both drift and diffusion coefficients depend on random and fixed effects simultaneously, presenting a novel generalization. We propose a novel parameter estimation method and provide theoretical insights into the asymptotic behavior of these estimators. Our estimation method diverges from the quasi-likelihood approach, offering a more accessible numerical procedure. This trade-off ensures stability and ease of implementation in high-frequency frameworks.
@@
The rest of this paper is organized as follows. Section 2 describes the precise model setup and assumptions. Section 3 introduces the stepwise estimation strategy with two kinds of the quasi-(log-)likelihood functions, and presents asymptotic properties of the associated estimators. In Section 4, we present numerical experiments illustrating the theoretical results, followed in Section 5 by some empirical applications.
@@ 2.1. Underlying dynamics
  • The driving noise processes $w_{1}(\cdot),w_{2}(\cdot),\dots$ are i.i.d. standard one-dimensional Wiener processes independent of the i.i.d. initial random variables $Y_{1}(0),Y_{2}(0),\dots$ in $\mathbb{R}$. We assume that for every $K>0$,
@@
      $\varphi_{r,i}\sim N_{p_{\varphi_{r}}}(\mu_{r},\Sigma_{r})$ (2.4)
      for $\mu_{r}\in\mathbb{R}^{p_{\varphi_{r}}}$ and positive definite $\Sigma_{r}\in\mathbb{R}^{p_{\varphi_{r}}}\otimes\mathbb{R}^{p_{\varphi_{r}}}$.
@@
The random-effect parameter $\tau_{i}$ represents the model-time scale of the $i$th individual; see Remark 3.1. Given the individual model-time scales, the random-effect parameter $\varphi_{r,i}$ represents variability in trend characteristics across $N$ individuals. For convenience, we denote ($p_{\varphi}:=p_{\varphi_{f}}+p_{\varphi_{r}}$)
$\varphi_{i}=(\varphi_{f},\varphi_{r,i})\in\mathbb{R}^{p_{\varphi}},\quad\mu=(\varphi_{f},\mu_{r}),\quad\Sigma=\mathrm{diag}(O,\Sigma_{r}).$ (2.5)
The random variable $\varphi_{i}$ is truly a degenerate Gaussian.
@@
$\vartheta:=(\eta,\theta_{\tau},\mu,\Sigma_{r})\in\Theta_{\eta}\times\Theta_{\tau}\times\Theta_{\mu}\times\Theta_{\Sigma_{r}}\subset\mathbb{R}^{p_{\eta}}\times\mathbb{R}^{p_{\tau}}\times\mathbb{R}^{p_{\varphi}}\times\mathbb{R}^{p_{\varphi_{r}}\times p_{\varphi_{r}}}.$ (2.6)
We write $\Theta:=\Theta_{\eta}\times\Theta_{\tau}\times\Theta_{\mu}\times\Theta_{\Sigma_{r}}$ for the full-parameter space. We assume that $\Theta_{\eta}\times\Theta_{\tau}$, $\Theta_{\mu}$, and $\Theta_{\Sigma_{r}}$ are a bounded convex domain and that any $A\in\overline{\Theta_{\Sigma_{r}}}$ is positive definite. In particular, the following form of the Sobolev inequality holds (for example, [1, Theorem 1.4.2]): for any $\mathcal{C}^{1}(\Theta_{\eta}\times\Theta_{\tau})$-function $\psi$ and $\alpha>\dim(\Theta_{\eta}\times\Theta_{\tau})=p_{\eta}+p_{\tau}$, we have
$\sup_{(\eta,\theta_{\tau})}|\psi(\eta,\theta_{\tau})|^{\alpha}\leq C\int_{\Theta_{\tau}}\int_{\Theta_{\eta}}(|\psi(\eta,\theta_{\tau})|^{\alpha}+|\partial_{(\eta,\theta_{\tau})}\psi(\eta,\theta_{\tau})|^{\alpha})d\eta d\theta_{\tau}$ (2.7)
for some universal constant $C=C(\Theta_{\eta},\Theta_{\tau},\alpha)>0$. The true value of $\vartheta$ by $\vartheta_{0}:=(\eta_{0},\theta_{\tau,0},\mu_{0},\Sigma_{r,0})\in\Theta$ with $\mu_{0}=(\varphi_{f,0},\mu_{r,0})$ is assumed to exist.
In this paper, we assume that we know a priori which component of the random effect parameter $\varphi_{i}$ is truly random; if not, the statistical model exhibits a non-standard feature that the true value of the covariance matrix $\Sigma$ can lie on the boundary of the parameter space and we will not touch this model-selection problem here; in such case, the regularization will help, see, for example, [4], [15], and [27].
@@
$\left\{Y_{i}(t_{j}):\leavevmode\nobreak\ 1\leq i\leq N,\quad 1\leq j\leq n\right\},$
@@ 2.2. Basic notation
  • For a matrix $A$, we write $A^{\otimes 2}=AA^{\top}$, where $\top$ stands for the transposition, and $\lambda_{\min}(A)$ denotes the minimum eigenvalue of $A$.
@@
  • For any measurable function $\mathsf{f}$ on $\mathbb{R}^{q}\times\mathbb{R}\times\overline{\Theta}$, we write
@@
    with simply $\mathsf{f}_{i}(t)$ and $\mathsf{f}_{i,j-1}$, respectively, when $\vartheta=\vartheta_{0}$ or the function $\mathsf{f}$ does not depend on $\vartheta$; for example, $S_{i}(t;\eta):=S(Y_{i}(t);\eta)$, $S_{i}(t):=S_{i}(t;\eta_{0})$, and so on.
@@ 2.3. Assumptions
We write
$\sup_{i}=\sup_{N\geq 1}\max_{i\leq N},\qquad\sup_{j}=\sup_{n\geq 1}\max_{j\leq n},\qquad\sup_{i,j}=\sup_{n,N\geq 1}\max_{i\leq N,\,j\leq n},$ (2.13)
where the supremum $\sup_{n,N\geq 1}$ is taken under (2.11).
@@
Note that the random elements $(\tau_{1},\varphi_{1},Y_{1}(\cdot)),(\tau_{2},\varphi_{2},Y_{2}(\cdot)),\dots$ are i.i.d. As in [6, Lemma 1], by Itô’s formula with the moment conditions (2.2) and (2.18) below, and also by the boundedness of $(a,b,c)$ and the Burkholder inequality, it is easy to show that for every $K>0$,
$\sup_{t\in[0,T]}E\big{[}|Y_{1}(t)|^{K}\big{]}+\sup_{j}\sup_{t\in(t_{j-1},t_{j}]}E\left[\left|\frac{1}{\sqrt{h}}(Y_{1}(t)-Y_{1}(t_{i,j-1}))\right|^{K}\right]<\infty.$ (2.16)
@@
  3. (3)
    The matrix ${\int_{0}^{T}\frac{a_{i}^{\otimes 2}(t)}{S_{i}(t)}dt}$ is a.s. positive definite.
@@
  1. (1)
    We assumed the boundedness of the coefficients in Assumption 2.1. What is essential in our technical proofs are the estimates (2.16). As was mentioned in [6, p.1932], if the support of $\mathcal{L}(\tau_{1},\varphi_{1})$ in $\mathbb{R}^{1+p_{\varphi}}$ is essentially bounded, then it is possible to relax (1) and (3) in Assumption 2.1 so that the coefficient $a(y)$ is at most polynomial growth. The reason is that the unboundedness of the support prevents us from using the standard Gronwall-inequality argument to deduce the moment boundedness of $E[|Y_{i}(t)|^{K}]$ and $E[|y_{ij}|^{K}]$, which is crucial in the proofs of the main results.
  2. (2)
    Assumption 2.2(2), which says that the function $x\mapsto\partial_{\eta}\log c(y;\eta)$ is not identically zero, is required to identify $\eta$ in the presence of the random-effect parameter $\tau_{i}$; for example, for a known and non-constant function $(c_{0},c_{1})$, this includes $c(y;\eta)=\exp\{c_{0}(y)+\eta\cdot c_{1}(y)\}$ while excludes the multiplicative case $c(y;\eta)=\eta c_{0}(y)$. The arguments are in line with those in [13], while there is some additional complexity here due to the presence of the random-effect parameters.
  3. (3)
    It is straightforward to extend the model (2.1) for $Y_{i}$ to a multivariate one, say in $\mathbb{R}^{d}$, where $a(y)$ is $\mathbb{R}^{d}\otimes\mathbb{R}^{p_{\varphi}}$-valued and $c(y;\eta_{1})$ is $\mathbb{R}^{d}\otimes\mathbb{R}^{r}$-valued and where $w_{i}$ is an $r$-dimensional standard Wiener process.
@@
where moreover, $f(\tau;\theta_{\tau})=f(\tau;\theta_{\tau,0})$ ($\mathcal{L}(\tau_{1})$-a.e.) if and only if $\theta_{\tau}=\theta_{\tau,0}$. Moreover,
$\forall\theta_{\tau}\in\overline{\Theta_{\tau}}\leavevmode\nobreak\ \forall K_{\tau}\in\mathbb{R},\quad\int_{0+}^{\infty}y^{K_{\tau}}f(y;\theta_{\tau})dy<\infty.$ (2.18)
@@ 3. Profile-marginal quasi-likelihood inference
In the estimation of a single diffusion-type model, it is possible to conduct a stepwise inference through profiling under a sufficiently high-frequency sampling scheme. It first focuses on the diffusion coefficient while ignoring the drift term, often enabling us to conduct an adaptive inference by reducing the computational costs in (numerical) optimizations: among others, we refer to [25] and [20] for details.
In the mixed-effects framework, the previous study [16, Section 3] developed a simple three-stage conditional least-squares estimation procedure in a class of generalized-hyperbolic type location-scale mixed-effects models. However, it will be impossible to directly apply the same strategy as in [16] to our model (2.1) since each conditional distribution of $Y_{i}$ given $(\varphi_{i},\tau_{i})$ is dynamic so that the filtration structure is much more complicated to handle. We will further modify the stepwise estimation procedure by profiling the random-effect parameters to obtain computationally convenient estimators; the strategy is possible due to the specific form of the coefficients of our model setup. Relatedly, [17, Section 4] studied the estimation of the distribution of the random-effect parameter in the drift coefficient based on continuous records. There, plug-in-type parametric estimation was considered for, as a specific example, the positive random-effect parameter $\varphi_{i}\sim\Gamma(a,\lambda)$ in $dY_{i}(t)=-\varphi_{i}Y_{i}(t)dt+dW_{i}(t)$ based on continuous records $\{(Y_{i}(t)_{t\leq T})\}_{i\leq N}$ for large $T$ and $N$. This work is an important theoretical study. However, in our discrete-time sampling situation, we need essentially different technical stuff.

### v2

@@ 1. Introduction
Mixed-effects models play a pivotal role in various scientific domains, facilitating the precise analysis of repeated observations among individuals. Despite their utility, parameter estimation in these models presents significant technical challenges. However, their application improves our understanding of dynamic biological phenomena, making them invaluable tools in numerous fields. Moreover, there is a growing interest in employing mixed-effects models coupled with stochastic differential equations (SDE). Originating from efforts in pharmacokinetics (see , e.g. references in [donnet2013review]), where the aim was to refine the modeling of intrinsic variability in drug distribution within organisms, these models have now branched out into diverse applications within life sciences. From the intricate workings of neurobiology ([ditlevsen2013introduction]) to the intricate dance of epidemic dynamics prediction ([narci2022inference]), mixed-effects models offer invaluable insight. Tailored estimation methodologies have been developed to fully exploit the potential of these models. Although numerous algorithms are proposed in the literature (see e.g. references in [picchiniWeb]), addressing parameter estimation in these models from a theoretical point of view remains rare. Indeed, maximum likelihood estimation, a favored method, poses challenges due to its rare explicit form. Only in select works do the authors provide specific results. Ditlevsen and De Gaetano are mainly limited to Brownian motion with random drift [ditlevsen2005mixed]. Then Delattre, Samson, Genon-Catalot, and Larédo focused on real diffusions with specific characteristics and Gaussian and inverse gamma random effects in the drift and in the diffusion coefficient, respectively ([DelGenLar18-2], [delattre2018parametric], [delattre2013maximum], [delattre2015estimation]). Covariates and Gaussian random effects in the drift were considered in [GroSamDit20] in multivariate diffusion models. However, there still remains a gap in understanding the asymptotic behavior of parameter estimators in more general mixed-effects diffusion models.
@@
Furthermore, prevalent assumptions in mixed-effects model literature, such as random effects following Gaussian distributions, may not always hold in practice. Recent advancements propose extensions, introducing distributions like the generalized hyperbolic distribution to better accommodate variability (see e.g. [AsaBolDigWal20] and [FujMas23]).
@@
Our work revisits the notion of utilizing normal variance-mean mixture-type random effects in broader mixed-effects models defined by stochastic differential equations. This allows for the exploration of a wider class of mixed-effects diffusion models compared to the previous literature. We also consider that both drift and diffusion coefficients depend on random and fixed effects simultaneously, presenting a novel generalization. We propose a novel parameter estimation method and provide theoretical insights into the asymptotic behavior of these estimators. Our estimation method diverges from the quasi-likelihood approach, offering a more accessible numerical procedure. This trade-off ensures stability and ease of implementation in high-frequency frameworks.
@@
The rest of this paper is organized as follows. Section 2 describes the precise model setup and assumptions. Section 3 introduces the stepwise estimation strategy with two kinds of quasi-(log-)likelihood functions, and presents asymptotic properties of the associated estimators. In Section 4, we present numerical experiments illustrating the theoretical results, followed in Section 5 by some empirical applications. Section 6 provides some discussion. Finally, Section A presents the proofs of the main result.
@@ 2.1. Underlying dynamics
  • The driving noise processes $w_{1}(\cdot),w_{2}(\cdot),\dots$ are i.i.d. standard one dimensional Wiener processes independent of the i.i.d. initial random variables $Y_{1}(0)$, $Y_{2}(0)$, $\dots$ in $\mathbb{R}$. We assume that for every $K>0$,
@@
      $\varphi_{r,i}\sim N_{p_{\varphi_{r}}}(0,\Sigma_{r})$ (2.4)
      for a positive definite $\Sigma_{r}\in\mathbb{R}^{p_{\varphi_{r}}}\otimes\mathbb{R}^{p_{\varphi_{r}}}$.
@@
The random-effect parameter $\tau_{i}$ represents the model-time scale of the $i$th individual relative to the common period $[0,T]$. How to incorporate the random-effect parameter $\tau_{i}$ in the drift part could be many things: see Remarks 3.1 and 3.7. Given the individual model-time scales, the random-effect parameter $\varphi_{r,i}$ represents variability in trend characteristics across $N$ individuals. Let $p_{\varphi}:=p_{\varphi_{f}}+p_{\varphi_{r}}$. To compactify the notation, we denote
$\begin{split}&\varphi_{i}=(\varphi_{f},\varphi_{r,i})\in\mathbb{R}^{p_{\varphi}},\\ &\mu=(\varphi_{f},0)\in\mathbb{R}^{p_{\varphi}},\quad\Sigma=\mathrm{diag}(O,\Sigma_{r})\in\mathbb{R}^{p_{\varphi}}\otimes\mathbb{R}^{p_{\varphi}}.\end{split}$ (2.5)
That is, we assume that we know a priori which component of the random effect parameter $\varphi_{i}$ is truly random; if not, the statistical model exhibits a non-standard feature that the true value of the covariance matrix $\Sigma$ can lie on the boundary of the parameter space and we will not touch this model-selection problem here; in such case, the regularization will help, see, for example, [BonKriGho10], [FanLi12], and [YosYos24]. We note, however, that our setting is implicitly able to consider the cases of non-degenerate $\varphi_{i}$ when $a_{f}(y)=a_{r}(y)$ (hence $p_{\varphi_{f}}=p_{\varphi_{r}}$); in this case, we may conveniently read $\varphi_{i}\sim N_{p_{\varphi_{f}}}(\varphi_{f},\Sigma_{r})$.
@@
$\vartheta:=(\eta,\theta_{\tau},\varphi_{f},\Sigma_{r})$ $\in\Theta_{\eta}\times\Theta_{\tau}\times\Theta_{\varphi_{f}}\times\Theta_{\Sigma_{r}}$
$\subset\mathbb{R}^{p_{\eta}}\times\mathbb{R}^{p_{\tau}}\times\mathbb{R}^{p_{\varphi_{f}}}\times\mathbb{R}^{p_{\varphi_{r}}\times p_{\varphi_{r}}}.$ (2.6)
We write $\Theta:=\Theta_{\eta}\times\Theta_{\tau}\times\Theta_{\varphi_{f}}\times\Theta_{\Sigma_{r}}$ for the full-parameter space. We assume that $\Theta_{\eta}\times\Theta_{\tau}$, $\Theta_{\varphi_{f}}$, and $\Theta_{\Sigma_{r}}$ are a bounded convex domain and that any $A\in\overline{\Theta_{\Sigma_{r}}}$ is positive definite. We denote the true value of $\vartheta$ by
$\vartheta_{0}:=(\eta_{0},\theta_{\tau,0},\varphi_{f,0},\Sigma_{r,0})\in\Theta,$ (2.7)
assumed to exist; also, we will write $\mu_{0}=(\varphi_{f,0},0)$ and $\Sigma_{0}=\mathrm{diag}(O,\Sigma_{r,0})$.
@@
$\left\{Y_{i}(t_{j}):\penalty 10000\ 1\leq i\leq N,\quad 1\leq j\leq n\right\},$
@@ 2.2. Basic notation
  • For a matrix $A$, we write $A^{\otimes 2}=AA^{\top}$, where $\top$ stands for the transposition, and $\lambda_{\min}(A)$ denotes the minimum eigenvalue of $A$ if it is square.
@@
  • For any measurable function $\mathsf{f}$ on $\mathbb{R}\times\overline{\Theta}$, we write
@@
    with simply $\mathsf{f}_{i}(t)$ and $\mathsf{f}_{i,j-1}$, respectively, when $\vartheta=\vartheta_{0}$ or the function $\mathsf{f}$ does not depend on $\vartheta$; for example, $S_{i}(t;\eta):=S(Y_{i}(t);\eta)$, $S_{i}(t):=S_{i}(t;\eta_{0})$, $S_{i,j-1}:=S_{i}(t_{j-1})$, and so on.
@@ 2.3. Assumptions
Note that the random elements $(\tau_{1},\varphi_{1},Y_{1}(\cdot)),(\tau_{2},\varphi_{2},Y_{2}(\cdot)),\dots$ are i.i.d. As in [delattre2018parametric, Lemma 1], by Itô’s formula with the moment conditions (2.2) and (2.17) below, and also by the boundedness of $(a,b,c)$ and the Burkholder inequality, it is not difficult to verify that for every $K>0$,
$\sup_{t\in[0,T]}E\big[|Y_{1}(t)|^{K}\big]+\sup_{n\geq 1}\max_{j\leq n}\sup_{t\in(t_{j-1},t_{j}]}E\left[\left|\frac{1}{\sqrt{h}}(Y_{1}(t)-Y_{1}(t_{j-1}))\right|^{K}\right]<\infty.$ (2.15)
@@
  3. (iii)
    The matrix ${\int_{0}^{T}\frac{a_{i}^{\otimes 2}(t)}{S_{i}(t)}dt}$ $(i\geq 1)$ is a.s. positive definite.
@@
  1. (i)
    We assumed the boundedness of the coefficients in Assumption 2.1. What is essential in our technical proofs are the estimates (2.15). As was mentioned in [delattre2018parametric, p.1932], if the support of $\mathcal{L}(\tau_{1},\varphi_{1})$ in $\mathbb{R}^{1+p_{\varphi}}$ is essentially bounded, then it is possible to relax (1) and (3) in Assumption 2.1 so that the coefficient $a(y)$ is at most polynomial growth. The reason is that the unboundedness of the support prevents us from using the standard Gronwall-inequality argument to deduce the moment boundedness of $E[|Y_{i}(t)|^{K}]$ and $E[|\overline{y}_{ij}|^{K}]$, which is crucial in the proofs of the main results.
  2. (ii)
    Assumption 2.2(2), which says that the function $x\mapsto\partial_{\eta}\log c(y;\eta)$ is not identically zero, is required to identify $\eta$ in the presence of the random-effect parameter $\tau_{i}$; for example, for a known and non-constant function $(c_{0},c_{1})$, this includes $c(y;\eta)=\exp\{c_{0}(y)+\eta\cdot c_{1}(y)\}$ but excludes the multiplicative case $c(y;\eta)=\eta c_{0}(y)$. The arguments are in line with those in [EguMas19], while there is some additional complexity here due to the presence of the random-effect parameters.
  3. (iii)
    It is straightforward to extend the model (2.1) for $Y_{i}$ to a multivariate one, say in $\mathbb{R}^{d}$, where $a(y)$ is $\mathbb{R}^{d}\otimes\mathbb{R}^{p_{\varphi}}$-valued and $c(y;\eta)$ is $\mathbb{R}^{d}\otimes\mathbb{R}^{r}$-valued and where $w_{i}$ is an $r$-dimensional standard Wiener process.
@@
and $f(\tau;\theta_{\tau})=f(\tau;\theta_{\tau,0})$ ($\mathcal{L}(\tau_{1})$-a.e.) if and only if $\theta_{\tau}=\theta_{\tau,0}$. Moreover,
$\forall K_{\tau}\in\mathbb{R},\quad\sup_{\theta_{\tau}\in\overline{\Theta_{\tau}}}\int_{0+}^{\infty}y^{K_{\tau}}f(y;\theta_{\tau})dy<\infty.$ (2.17)
@@ 3. Profile-marginal quasi-likelihood inference
In the estimation of a single diffusion-type model, it is possible to conduct a stepwise inference through profiling under a sufficiently high-frequency sampling scheme. It first focuses on the diffusion coefficient while ignoring the drift term, often enabling us to conduct an adaptive inference by reducing the computational costs in (numerical) optimizations: among others, we refer to [UchYos12] and [KamUch15] for details.
In the mixed-effects framework, the previous study [FujMas23, Section 3] developed a simple three-stage conditional least-squares estimation procedure in a class of generalized hyperbolic type location-scale mixed-effects models. However, it will be impossible to directly apply the same strategy as in [FujMas23] to our model (2.1) since each conditional distribution of $Y_{i}$ given $(\varphi_{i},\tau_{i})$ is dynamic, so that the filtration structure is much more complicated to handle. We will further modify the stepwise estimation procedure by profiling the random-effect parameters to obtain computationally convenient estimators; the strategy is possible due to the specific form of the coefficients of our model setup. Relatedly, [GenLar16, Section 4] studied the estimation of the distribution of the random-effect parameter in the drift coefficient based on continuous records. There, plug-in-type parametric estimation was considered for, as a specific example, the positive random-effect parameter $\varphi_{i}\sim\Gamma(a,\lambda)$ in $dY_{i}(t)=-\varphi_{i}Y_{i}(t)dt+dW_{i}(t)$ based on continuous records $\{(Y_{i}(t))_{t\leq T})\}_{i\leq N}$ for large $T$ and $N$. This work is an important theoretical study. However, in our discrete-time sampling situation, we need essentially different technical stuff.